a company gives its employees a 4% raise at the beginning of every year.This year Jordan earns $32000. Write a rule to find Jordan's salary after n years?
Also, how many years will it take Jordan to earn $40,000?

Answer by Boreal(15235)

The general rule is his salary will be 32000*(1.04)^n, where n is the number of years.
40,000=32,000(1.04)^n
divide both sides by 32,000
(5/4)=1.04^n
take ln both sides
ln(5/4)=nln(1.04)
divide ln (5/4)/ln(1.04)=5.69 years, so it will take 6 years to exceed $40,000. $32,000*(1.04)^6=$40,490